OpenStudy (nsh4267):

A contractor finds that it takes Sam Tockett 9 hours to construct a wall of a certain size. It takes Jim Lowell 7 hours to contruct the same wall. How long would it take if they worked together?

OpenStudy (anonymous):

\[\frac{9\times 7}{9+7}\] then arithmetic

OpenStudy (anonymous):

\[\frac{63}{16}=3\tfrac{15}{16}\]

OpenStudy (nsh4267):

so about 3.9 hours

OpenStudy (anonymous):

we can work out the reasoning if you like one rate is \[\frac{1}{7}\] per hour, other rate is \[\frac{1}{9}\] per hour, combined rate is \[\frac{1}{7}+\frac{1}{9}=\frac{7+9}{7\times 9}\] and you want to solve for time in \[\frac{7+9}{7\times 9}T=1\] making \[T=\frac{7\times 9}{7+9}\]

OpenStudy (nsh4267):

I think I have it figured out but I got 3.9 hours working together.

OpenStudy (anonymous):

Total = 63/ 16 = 4 hrs